### v1

# Differentiable Energy-Based Regularization in GANs: A Simulator-Based Exploration of VQE-Inspired Auxiliary Losses

Abstract
This paper presents an exploratory, simulator-based proof of concept investigating whether differentiable energy terms derived from parameterized quantum circuits can serve as auxiliary regularization signals in Generative Adversarial Networks (GANs). We augment the Auxiliary Classifier GAN (ACGAN) generator objective with a Variational Quantum Eigensolver (VQE)-inspired energy term computed from class-specific Ising Hamiltonians using Qiskit’s EstimatorQNN and TorchConnector.
[…]
On MNIST, we observe that the energy-regularized model (termed QACGAN) achieves high classification accuracy (99–100%) within 5 epochs compared to 87.8% for ACGAN, suggesting the auxiliary term influences class conditioning. However, sample quality metrics (FID) show high variance across runs (coefficient of variation $\approx$25% at epoch 5), with values ranging from 19.92 to 35.96. Extended runs stabilize around FID 23–24, comparable to the ACGAN baseline.
We explicitly do not claim quantum advantage, improved stability in any general sense, or scalability beyond this toy setting. The contribution is methodological: demonstrating that VQE-style energy computations can be integrated into GAN training loops via differentiable pathways. Whether such auxiliary signals provide benefits beyond equivalent classical regularizers remains an open question requiring systematic ablation studies, which we leave for future work.
Keywords: Energy-Based Regularization, Generative Adversarial Networks, Variational Quantum Eigensolver, Hybrid Quantum-Classical Algorithms, Proof of Concept, ACGAN.
1 Introduction
Generative Adversarial Networks (GANs) [goodfellow2014generative] have achieved remarkable success in image synthesis, but remain prone to mode collapse and unstable dynamics, e.g., vanishing or exploding gradients [salimans2016improved, arjovsky2017wasserstein, gulrajani2017improved], see also broader overviews [purwono2025understanding, kossale2022mode]. The Auxiliary Classifier GAN (ACGAN) [odena2017conditional] adds a class-prediction head to the discriminator, often improving class-conditional fidelity. Yet fundamental training issues can persist.
In parallel, the Noisy Intermediate-Scale Quantum (NISQ) era [preskill2018quantum] has motivated hybrid schemes that combine quantum circuits with classical optimizers. The Variational Quantum Eigensolver (VQE) [peruzzo2014variational, mcclean2016theory] is a leading approach for estimating ground-state energies of Hamiltonians using parameterized circuits and classical optimization.
Scope and intent of this work. This paper is an exploratory proof of concept investigating whether VQE-style energy computations can be integrated into GAN training as differentiable auxiliary losses. We do not claim that the quantum aspect provides any advantage over classical alternatives—indeed, we explicitly lack ablation studies comparing against equivalent classical biases. Our contribution is purely methodological: demonstrating the technical feasibility of backpropagating gradients through a VQE energy term into a GAN generator.
We explore a hybrid architecture, QACGAN, that supplements the generator objective with a VQE-based energy term computed from a class-specific Ising Hamiltonian. The Hamiltonian parameterization is intentionally simple and serves only to induce class-dependent separation rather than encode meaningful physics. In our present implementation, the energy is computed on a noiseless statevector simulator and integrated via a differentiable pathway (EstimatorQNN + TorchConnector), allowing gradients to backpropagate through the quantum module.
2 Background
2.1 Auxiliary Classifier GAN (ACGAN)
ACGAN [odena2017conditional] extends conditional GANs by having the discriminator $D$ predict both source (real/fake) and class labels, while the generator $G$ consumes noise $z$ and class $c$ to synthesize $X_{\text{fake}}=G(z,c)$. The discriminator maximizes $\mathcal{L}_{S}+\mathcal{L}_{C}$ (source and class log-likelihoods). In the original formulation, the generator maximizes $\mathcal{L}_{C}-\mathcal{L}_{S}$; in our implementation (and most practical implementations), both objectives are cast as minimization problems via negation and standard loss functions (BCE, cross-entropy), so the generator effectively minimizes a combined adversarial and auxiliary loss.
2.2 Variational Quantum Eigensolver (VQE)
Given a Hamiltonian $\mathcal{H}=\sum_{i}c_{i}P_{i}$ (Pauli decomposition), VQE minimizes the energy $\langle\psi(\vec{\theta})|\mathcal{H}|\psi(\vec{\theta})\rangle\geq E_{G}$ via a parameterized ansatz $|\psi(\vec{\theta})\rangle=U(\vec{\theta})|0\rangle^{\otimes n}$ and a classical optimizer [peruzzo2014variational, mcclean2016theory].
3 Method: QACGAN with VQE Regularization
We associate each class $c$ with an $N$-qubit Ising Hamiltonian
$H_{c}=-J\sum_{\langle i,j\rangle}\sigma_{z}^{i}\sigma_{z}^{j}-\sum_{i=1}^{N}h_{c,i}\,\sigma_{z}^{i},$ (1)
where $J$ is a fixed coupling coefficient (set to $J=1.0$ in our experiments) and $h_{c,i}$ are class-specific local fields [lucas2014ising]. The negative signs correspond to a ferromagnetic Ising model where lower energies favor aligned spins. In our experiments, we use $N=4$ qubits and a simple nearest-neighbor (linear chain) coupling topology. We emphasize that this parameterization is intentionally minimal and primarily serves to induce a class-dependent separation in the energy term, rather than to encode a physically meaningful model of the data.
Class-specific local fields.
In our implementation, the class-specific local fields $h_{c,i}$ are initialized deterministically using a simple linear rule. For each class $c\in\{0,\dots,9\}$, all qubits share the same field value
$h_{c,i}\;=\;h_{\text{global}}+c\cdot\Delta h\qquad\forall i\in\{1,\dots,N\},$ (2)
where $h_{\text{global}}=0.1$ and $\Delta h=0.01$. Thus, the fields span the range $[0.10,0.19]$ across classes. The Hamiltonians are constructed once during initialization and remain fixed throughout training; no learning or adaptation of $h_{c,i}$ is performed. The role of the Hamiltonian in this work is therefore not to model domain-specific physics, but to provide a simple, class-dependent energy landscape acting as an auxiliary regularization signal in the generator objective. A small MLP (AngleProducer) maps the element-wise product of noise $z$ and class embedding to circuit parameters $\vec{\theta}(z,c)$ for a Qiskit EfficientSU2 ansatz (4 qubits, one repetition) [kandala2017hardware]. The energy
$E_{c}(z)=\big\langle\psi(\vec{\theta}(z,c))\big|H_{c}\big|\psi(\vec{\theta}(z,c))\big\rangle$ (3)
is evaluated with Qiskit’s StatevectorEstimator (noiseless statevector simulation).
Generator objective.
We add a VQE penalty to the ACGAN generator loss:
$\mathcal{L}_{G}\;=\;\mathcal{L}_{\text{adv}}+\mathcal{L}_{\text{aux}}\;+\;\lambda_{\text{VQE}}\,\mathcal{L}_{\text{VQE}},\qquad\text{with }\mathcal{L}_{\text{VQE}}:=E_{c},$ (4)
where $\mathcal{L}_{\text{adv}}$ is the adversarial loss (binary cross-entropy targeting the "real" label) and $\mathcal{L}_{\text{aux}}$ is the auxiliary classification loss (cross-entropy for class labels).
Implementation note. In this work, $\mathcal{L}_{\text{VQE}}$ is implemented using Qiskit’s EstimatorQNN with TorchConnector, ensuring that gradients from the VQE energy evaluation propagate through the quantum circuit into the AngleProducer and ultimately into the generator. Energies are computed on a noiseless StatevectorEstimator backend [qiskit, qiskit_ml_docs].
Related work.
Quantum generative modeling has been explored along several lines. Dallaire-Demers and Killoran proposed a quantum–classical adversarial setup where a parameterized quantum circuit plays the role of a generator trained against a classical discriminator [dallaire2018qgan]. Benedetti et al. studied adversarial quantum circuit learning for pure state approximation, highlighting training dynamics and optimization aspects in fully or partially quantum settings [benedetti2019aqc]. Zoufal et al. introduced QGANs for loading classical probability distributions into quantum states, demonstrating gradient-based training with a classical discriminator [zoufal2019qgan]. In contrast to these lines, our approach does not pursue an adversarial quantum generator end-to-end; instead, we inject a differentiable VQE energy as a class-specific prior directly into the classical ACGAN generator objective. This hybrid objective provides an interpretable regularization signal that stabilizes early training while keeping the main image-synthesis pipeline classical.
4 Experiments
4.1 Setup
Data. MNIST [lecun1998gradient].
Architectures. $G$ and $D$ follow DCGAN-style backbones [radford2015unsupervised]. In our implementation, both ACGAN and QACGAN use heads realized by flattening $4{\times}4$ features followed by fully connected layers (no $1{\times}1$ convolutional heads).
Optimization. We use Adam with learning rate $2{\times}10^{-4}$, $\beta_{1}{=}0.5$, $\beta_{2}{=}0.999$. The effective batch size of $64$ is achieved by gradient accumulation of two steps over micro-batches of size $32$. Real-label smoothing is set to $0.9$ [salimans2016improved]. We use fixed random seeds for reproducibility: the 5-epoch QACGAN exploratory run and the first 10-epoch run use seed $42$, while the second 10-epoch run uses seed $2025$ to assess robustness across initializations. ACGAN is trained for 50 epochs. QACGAN is primarily evaluated after 5 epochs (due to simulator cost), with additional runs extended to 10 epochs for stability analysis. For QACGAN we set $\lambda_{\text{VQE}}=0.1$ and employ the Qiskit EfficientSU2 ansatz (repetitions =1=1, circular entanglement) on 4 qubits.
Quantum simulation. Energies are computed with Qiskit’s StatevectorEstimator in noiseless mode. No hardware noise model or error mitigation is applied.
Evaluation. At each evaluation step we compute: (i) Fréchet Inception Distance (FID) [heusel2017gans] and Inception Score (IS) [salimans2016improved] on 1,000 generated samples (using torch-fidelity with an MNIST reference set [torchfidelity]), and (ii) classification accuracy of a pretrained external CNN on 500 generated samples. Wall-clock timing is measured on the same GPU host; QACGAN training is dominated by the quantum simulation. We compute FID/IS using MNIST-domain features to maintain domain compatibility with generated digits rather than ImageNet-based embeddings.
4.2 Results
Metric ACGAN (50 ep.) QACGAN (5 ep.) QACGAN (10 ep., run 1) QACGAN (10 ep., run 2)
Seed 42 42 2025
Best FID $\downarrow$ 24.02 (@20 ep.) 19.92 (@5 ep.) 23.91 (@10 ep.) 23.23 (@9 ep.)
FID @ epoch 5 19.92 35.96 24.11
Best IS $\uparrow$ 2.23 (@25 ep.) 2.07 (@5 ep.) 2.29 (@2 ep.) 2.32 (@4 ep.)
Accuracy @ 5 ep. 87.8% 99.0% 99.0% 100.0%
Total training time 0h 27m 42s 7h 17m 36s 14h 2m 17s 14h 9m 1s
Table 1: MNIST results. QACGAN runs differ in number of epochs and random seeds. The 5-epoch run achieved the best single FID of 19.92; however, the 10-epoch runs with different seeds show FID values of 35.96 and 24.11 at epoch 5, indicating notable variance across runs (coefficient of variation $\approx$25% at epoch 5). Best overall FID across extended runs stabilizes around 23–24. All quantum runs use a noiseless StatevectorEstimator.
Figure 1: Inception Score across epochs for all runs. QACGAN achieves competitive or superior scores in early epochs, with peak values above ACGAN, though with greater variance.
Figure 2: FID across epochs for all runs. QACGAN reaches its best FID early (epoch 5) and remains competitive, while ACGAN requires more epochs and eventually exhibits degradation in sample quality.
Figure 3: External-classifier accuracy on generated samples across all runs. QACGAN models reach near-perfect class consistency within the first 5 epochs, surpassing the baseline ACGAN, which converges more slowly.
Figure 4: Comparison of generated MNIST samples. Left: baseline ACGAN (best FID =24.02=24.02 at 20 epochs). Right: QACGAN from the 5-epoch exploratory run (FID =19.92=19.92). Note that this represents the best observed QACGAN result; other runs showed higher FID at epoch 5 (see Table 1). QACGAN digits appear sharper and more class-consistent, consistent with higher external-classifier accuracy.
Table 1 summarizes the outcomes. The ACGAN baseline reached its best FID of $24.02$ after 20 epochs, best IS of $2.23$ at epoch 25, and $87.8\%$ classifier accuracy at epoch 5. In contrast, QACGAN consistently matched or surpassed these baselines within significantly fewer epochs.
The initial 5-epoch exploratory run achieved the lowest observed FID of $19.92$ together with $99.0\%$ classifier accuracy. However, we note substantial variance across runs at epoch 5: the two 10-epoch runs achieved FID values of $35.96$ (run 1, seed 42) and $24.11$ (run 2, seed 2025) at the same epoch, yielding a coefficient of variation of approximately 25%. This variance highlights the sensitivity of early-epoch GAN metrics to initialization and underscores the need for multiple runs when reporting results.
Despite this variance at epoch 5, both extended runs converged to similar best FID values: $23.91$ (run 1, epoch 10) and $23.23$ (run 2, epoch 9). Importantly, these longer runs achieved the strongest Inception Scores, $2.29$ and $2.32$, and classifier accuracies of $99.0\%$ and 100.0% at epoch 5. These results suggest that while early-epoch FID can vary substantially, QACGAN provides consistent improvements in class consistency (accuracy) and eventually stabilizes to competitive FID values.
Qualitatively, QACGAN samples after 5 epochs display sharper digits and clearer class identity compared to ACGAN at the same epoch. Figure 4 illustrates this difference.
Compute cost.
One QACGAN epoch requires $\approx 1.45$ h on noiseless statevector simulation, implying 7788 h for 5 epochs and about $14$ h for 10 epochs, compared to only $\approx 22$ s per epoch for the ACGAN baseline. This overhead is dominated by repeated evaluations and gradient propagation through Qiskit’s EstimatorQNN via the TorchConnector, which incurs substantial Python-level overhead in our current implementation. It should therefore be interpreted as an implementation and simulator artifact rather than an inherent computational cost of the quantum model itself.
5 Discussion
5.1 Interpretation of Results
Our experiments demonstrate that integrating a VQE-based energy signal into the generator objective influences early-stage training dynamics of ACGAN on MNIST. The energy-regularized model achieves consistently high class consistency (99–100% accuracy) within the first few epochs, compared to 87.8% for the baseline. However, FID values show substantial variance across runs (19.92–35.96 at epoch 5), and the effect on sample quality remains inconclusive.
What we can conclude: The auxiliary energy term appears to influence class conditioning, as evidenced by consistently high accuracy across all runs. The mechanism likely involves the class-specific Hamiltonian providing an additional gradient signal that encourages class separation.
What we cannot conclude: We cannot claim that this represents a “quantum advantage” or that the effect could not be replicated with an equivalent classical bias term. Without ablation studies comparing against classical regularizers with similar structure, the observed effects may simply reflect the presence of any auxiliary class-dependent signal rather than anything specific to the VQE formulation.
5.2 Critical Limitations
Missing ablation against classical baselines. The most significant limitation of this work is the absence of comparison against classical alternatives. A fair evaluation would require testing against: (1) a classical MLP producing the same class-dependent scalar, (2) simple class-conditional biases, or (3) other auxiliary loss formulations. Without such comparisons, we cannot attribute the observed effects to the quantum/VQE aspect specifically.
Trivial Hamiltonian design. The Ising Hamiltonian uses a deliberately simple linear parameterization ($h_{c,i}=0.1+0.01c$) that encodes no meaningful physics or problem structure. This choice was intentional for this proof-of-concept, but it means the “quantum” component is essentially computing a class-dependent scalar through an unnecessarily expensive pathway.
[…]
High variance and limited runs. FID variance of $\approx$25% at epoch 5 across runs with the same hyperparameters indicates high sensitivity to initialization. With only 3–4 runs per configuration, we lack statistical power to make strong claims about the method’s reliability.
5.3 Scope of Claims
We explicitly frame this work as boundary exploration rather than a performance contribution:
  • We demonstrate technical feasibility of integrating VQE energy computations into GAN training via differentiable pathways.
  • We observe that class-dependent auxiliary losses can influence accuracy metrics, but do not claim this effect is specific to quantum formulations.
  • We do not claim quantum advantage, improved stability in any general sense, or scalability beyond this toy MNIST setting.
5.4 Future Work
The most critical next step is ablation studies comparing against equivalent classical biases. Additional directions include: testing on noise-aware backends, exploring more expressive Hamiltonian designs, extending to larger datasets, and conducting statistically rigorous multi-run evaluations with confidence intervals.
Broader Context
This study is best viewed as methodological groundwork rather than a performance contribution. The integration of differentiable energy-based terms into generative models is independently interesting, and VQE provides one pathway for computing such terms. Whether quantum formulations offer any advantage over classical alternatives in this context remains an open question that this work does not—and does not attempt to—answer.
6 Conclusion
We presented an exploratory proof of concept investigating VQE-inspired energy terms as auxiliary regularization signals in GANs. On MNIST, the energy-regularized model achieved high classification accuracy (99–100%) within five epochs, compared to 87.8% for the baseline. Sample quality metrics (FID) showed high variance across runs and converged to values comparable to the baseline in extended training.
Key takeaway: This work demonstrates the technical feasibility of integrating VQE-style computations into GAN training loops via differentiable pathways. However, we explicitly do not claim quantum advantage, and the absence of ablation studies against classical alternatives means we cannot attribute observed effects to the quantum formulation specifically.
The contribution is methodological: showing that such integration is possible and produces non-trivial training dynamics. Whether VQE-based regularization offers any benefit beyond equivalent classical approaches remains an open question for future systematic investigation.
Acknowledgments
The author would like to express sincere gratitude to Prof. Roman Šenkeřík for his supervision and support during the preparation of this work.